OpenStudy (anonymous):

Rewrite the given equation in slope-intercept form: 15x - 3y = 12

OpenStudy (anonymous):

Anyone??

OpenStudy (ash2326):

@smartrob This section is for feedback. Could you post this in Maths subject? I'll help you there !

OpenStudy (anonymous):

slope intercept form is y=mx+b, where m is the slope and b is the y intercept. So step by step converting 15x-3y=12 to slope intercept form: 15x-3y=12 15x-12=3y divide both sides by 3... 5x-4=y or **y=5x-4** the slope is 5, the y intercept is -4